# Fair Graph Machine Learning under Adversarial Missingness Processes

### v4

@@ Abstract
Graph Neural Networks (GNNs) have achieved state-of-the-art results in many relevant tasks where decisions might disproportionately impact specific communities. However, existing work on fair GNNs often assumes that either sensitive attributes are fully observed or they are missing completely at random. We show that an adversarial missingness process can inadvertently disguise a fair model through the imputation, leading the model to overestimate the fairness of its predictions. We address this challenge by proposing Better Fair than Sorry (BFtS), a fair missing data imputation model for sensitive attributes. The key principle behind BFtS is that imputations should approximate the worst-case scenario for fairness—i.e. when optimizing fairness is the hardest. We implement this idea using a 3-player adversarial scheme where two adversaries collaborate against a GNN classifier, and the classifier minimizes the maximum bias. Experiments using synthetic and real datasets show that BFtS often achieves a better fairness $\times$ accuracy trade-off than existing alternatives under an adversarial missingness process.
@@ 1 Introduction
With the increasing popularity of machine learning in high-stakes decision-making, it has become a consensus that these models carry implicit biases, and these biases should be addressed to improve the fairness of algorithmic decisions ghallab2019responsible . The disparate treatment of such models towards African Americans and women has been illustrated in the well-documented COMPAS b1 and Apple credit card apple cases, respectively. While there has been extensive research on fair machine learning, the proposed solutions have mostly disregarded important challenges that arise in real-world settings. For instance, in many applications, data can be naturally modeled as graphs (or networks), representing different objects, their relationships, and attributes instead of as sequences or images.
@@
Figure 1: Motivation: In Figure (a), a graph machine learning algorithm is applied to decide who receives credit (positive or negative) based on a possibly missing sensitive attribute, gender (and binary only for illustrative purposes). As shown in Figure (b), traditional missing data imputation does not account for outcomes (positive/negative), and thus, their imputed values can under-represent the bias of the complete dataset—demographic parity (DP) is $0.09$ in this example. This paper proposes BFtS, an imputation method for graph data that optimizes fairness in the worst-case imputation scenario using adversarial learning, as shown in Figure (c), where DP is $0.47$.
Another challenge for fair machine learning is the pervasiveness of missing data little2019statistical . This is particularly critical because fair algorithms often require knowledge of sensitive attributes that are more likely to be missing due to biases in the collection process or privacy concerns. As an example, the absence of data in census surveys and health surveys is correlated with sensitive attributes such as gender, age, race, etc. o2019potential ; weber2021gender . Missing data is also a concern in networked data, especially when tracking the transmission of diseases ghani1998sampling . As a result, many fair machine learning algorithms rely on imputation methods to handle missing data and mitigate associated biases. However, when the missingness mechanism is not random, standard imputation techniques can introduce systematic errors that propagate through the pipeline and compromise downstream fairness mansoor2022impact ; jeong2022fairness .
This work investigates the impact of adversarial sensitive value missingness processes on fairness, where the pattern of missing sensitive data is structured to obscure true disparities. If the predictions from the imputation method fail to capture the true distribution of a sensitive attribute, any fairness-aware model trained on the imputed data is prone to inheriting the underlying true bias. More specifically, a key challenge arises when an adversarial missingness process makes the imputed dataset appear to be fair. In graphs, an adversary can exploit the graph structure to manipulate the imputed values and, as a consequence, the fairness-aware model. Prior work on the fairness of graph machine learning assumes sensitive values are Missing Completely At Random (MCAR). However, this assumption rarely holds in practice o2019potential ; weber2021gender ; ghani1998sampling ; jeong2022fairness . Methods that overlook the adversarial missingness process risk producing misleading fairness guarantees w.r.t. the complete data.
@@
We summarize the contributions of this paper as follows: (1) We investigate theoretically and empirically the potential of an adversarial missingness process to bias a fair graph machine learning model; (2) We propose Better Fair than Sorry (BFtS), a novel 3-player adversarial learning framework for the imputation of missing sensitive data that produces worst-case imputed values for fair GNNs and is effective under adversarial missingness processes and even when sensitive attribute information is completely unavailable; and (3) We show empirically that BFtS achieves a better fairness $\times$ accuracy trade-off than the existing alternatives.
@@ 1.1 Related Work
Fairness in graph machine learning. We divide the work in this area into three main directions (pre-processing, extended objective functions, and adversarial learning). FairOT, FairDrop, FairSIN, EDITS, and Graphair are pre-processing techniques that remove the bias from the graph before training fairot ; fairdrop ; fairsin ; edits ; graphair . Fairwalk, Crosswalk, Debayes, MONET, NIFTY, FairVGNN, PFR-AX, FairSAD, and FairGAE modify the objective function of GNNs to learn fair representations during training fairwalk ; crosswalk ; debayes ; monet ; agarwal2021towards ; wang2022improving_short ; merchant2023disparity ; zhu2024fair ; fan2021fair . CFC, FLIP, DKGE, and Debias train adversarial networks along with GNN for fair prediction cfc ; cfc2 ; kg ; zhang2018mitigating . However, different from our work, all these methods require fully observed sensitive attributes to incorporate fairness considerations.
@@
Missing Data Imputation and fairness. Missing data can be imputed using unconditional mean, reconstruction, preferential attachment, autoencoders, etc. donders2006gentle ; pereira2020reviewing ; huisman2009imputation . The traditional procedure for handling missing data is independent imputation—i.e., to first impute the missing values and then solve the task rossi2021unreasonable ; buck1960method . SAT is a matching-based GNN for graphs with missing attributes chen2020learning , but it does not account for fairness. Training a classifier from imputed data can amplify the bias of a machine learning model, as discussed in subramoniandiscrimination ; zhang2021assessing ; feng2023adapting ; guha2024automated ; martinez2019fairness ; fricke2020missing ; mansoor2022impact . Some studies try to generate fair feature imputations subramoniandiscrimination ; feng2023adapting ; jeong2022fairness ; zhang2022fairness . However, as the approaches discussed earlier, these methods do not consider missing sensitive information and require full knowledge of the sensitive attributes for fairness intervention.
@@
Fairness with missing sensitive attributes. These studies are the most related to ours. One extreme assumption is the complete unavailability of sensitive information. For instance, hashimoto2018fairness takes into account the worst-case distribution across a given group size, and lahoti2020fairness adversarially reweighs training samples to optimize the average loss. In chai2022self , a top $k$ average loss minimization scheme is applied for fairness. In zhao2022towards , the correlation between the predictions and other features associated with sensitive attributes is minimized. Other approaches to achieve fairness are class re-balancing yan2020fair and chai2022fairness using soft labels from an overfitted teacher model. However, in practice, some sensitive information is often available and can significantly improve fairness chai2022self . The above-mentioned models fail to utilize the accessible sensitive values. FairGNN fairgnn assumes limited sensitive information and learns a fair GNN by predicting the missing values with an independent imputation model. FairAC guo2023fair assumes that both node features and the sensitive attribute are missing for a subset of nodes, and utilizes only the observed sensitive attributes without applying any imputation. RNF du2021fairness generates proxy annotations for missing attributes using generalized cross-entropy zhang2018generalized . Similar to RNF, our approach (BFtS) handles completely or partially missing sensitive information. BFtS applies a 3-player scheme to minimize the maximum possible bias and outperforms FairGNN and RNF in terms of fairness $\times$ accuracy. Minimizing the maximum instead of the average risk, similar to our method, has been shown to achieve better guarantees shalev2016minimizing . A similar minmax approach has been applied to maximize the robustness and accuracy of uncertainty models lofberg2003minimax ; lanckriet2002robust ; chen2017robust ; fauss2021minimax . In nguyen2017dual ; vandenhende2019three , a 3-player adversarial network is proposed to improve the classification and stability of adversarial learning.
Adversarial attacks on fairness. Recent work has investigated data poisoning as a means to adversarially degrade model fairness mehrabi2021exacerbating ; solans2020poisoning . Some approaches, such as UnfairTrojan and TrojFair introduce backdoor attacks specifically aimed at reducing model fairness furth2024unfair ; zheng2023trojfair . In the context of graph neural networks (GNNs), adversarial fairness attacks have predominantly focused on modifying graph topology. For instance, hussain2022adversarial ; zhang2023adversarial alter the structure of the graph to compromise the fairness of GNN predictions. NIFA luo2024your attacks GNN fairness by injecting new nodes guided by an uncertainty-maximization and homophily-enhancement strategy. Similarly, FATE kang2023deceptive employs a bilevel optimization-based meta-learning framework to carry out poisoning attacks on GNNs. These existing approaches primarily target fairness through data poisoning, node injection, or structural perturbations. To the best of our knowledge, no prior work has explored attacks that can only manipulate the missingness process for sensitive attributes as a means to compromise model fairness.
@@ 2 Preliminaries
In this section, we formalize the setting considered in this paper and introduce the Graph Neural Network model, which is our base model for node classification. For simplicity, we only consider binary-sensitive attributes. By modifying the loss function in Eq. 2 to categorical cross entropy and MSE, our work can be generalized for non-binary and continuous sensitive attributes, respectively.
2.1 Notation and Problem Definition
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X},\mathcal{S})$ be an undirected graph where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, $\mathcal{X}$ are node attributes, and $\mathcal{S}$ is the set of sensitive attributes. The matrix $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix of $\mathcal{G}$ where $A_{uv}=1$ if there is an edge between $u$ and $v$ and $A_{uv}=0$, otherwise. We focus on the setting where sensitive attributes might be missing and only $\mathcal{V}_{S}\subseteq\mathcal{V}$ nodes include the information of the sensitive attribute $s_{v}$ for a node $v$. The sensitive attribute forms two groups, which are often called sensitive ($s_{v}=1$) and non-sensitive ($s_{v}=0$).
We consider the fair node classification task, where the goal is to learn a classifier $f_{C}$ to predict node labels $y_{v}\in\{0,1\}$ based on a training set $\mathcal{V}_{L}\subseteq\mathcal{V}$. Without loss of generality, we assume that the class $y=1$ is the desired one (e.g., receive credit or bail). Given a classification loss $\mathcal{L}_{C}$ and a fairness loss $\mathcal{L}_{A}$, the goal is to minimize their combination:
$f_{C}^{*}=\arg\min_{f_{C}}\mathcal{L}_{C}(f_{C})+\alpha\mathcal{L}_{A}(f_{C})$
Computing fairness loss $\mathcal{L}_{A}$ requires sensitive attributes $s_{v}$. The main challenge addressed in this paper is due to our assumption that sensitive attributes can be missing. In this scenario, one can apply an independent imputation model $s_{v}\approx f_{I}(v)$ to predict missing sensitive attributes before training the fair classifier. However, the fairness of the resulting classifier will be highly dependent on the accuracy of $f_{I}$, which is the key motivation for our work. Two of the models (or players) learned by our model are GNNs. The details about Graph Neural Networks are provided in Appendix A.
@@ 3 Introducing Bias via an Adversarial Missingness Process
Our work investigates the role of missing sensitive values in the fairness of graph machine learning. Datasets with prevalent missing values often require the application of a missing data imputation method before they can be processed by a fair machine learning model. In this section, we focus on how the missingness process (i.e., the process that generates missing values) can lead to (intended or unintended) biases in a fair model trained using the imputed data. More specifically, the missingness process can induce the fair model to underestimate the bias in the imputed data. As a consequence, the fair model applied to the imputed data can still be inadvertently biased w.r.t. the complete data.
The simplest missingness process for sensitive values is missing completely at random—i.e., the probability of a value being missing is independent of the data. However, in practical scenarios, missing values are rarely random o2019potential ; weber2021gender ; ghani1998sampling . We focus on the particular case where the missingness process is adversarial and unknown. We prove that manipulating the missingness process to induce bias optimally is computationally hard. However, we show empirically that a simple heuristic can effectively introduce bias to an independent imputation model that simply tries to maximize imputation accuracy. We formulate the problem for a given number of observed sensitive values $k$.
@@
Adversarial Missingness for Minimizing Classification Bias (AMMCB): Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with a feature matrix $\mathcal{X}$, hypothesis spaces $\Theta_{C}$ and $\Theta_{I}$ and a training dataset $\mathcal{V}_{L}$, select a set of nodes $\mathcal{V}_{S}^{*}\subset\mathcal{V}$ such that $|\mathcal{V}_{S}^{*}|=k$ as to minimize the fairness loss $\mathcal{L}_{A}$ of the optimal classifier $f_{C}^{*}\in\Theta_{C}$ when it is trained with sensitive values imputed from the optimal imputation model $f_{I}^{*}$ that itself is trained with $\mathcal{V}_{S}^{*}$.
AMMCB is a discrete tri-level optimization problem as $f_{I}^{*}$ depends on $\mathcal{V}_{S}^{*}$ and $f_{C}^{*}$ depends on $f_{I}^{*}$. We note that by minimizing the bias loss $\mathcal{L}_{A}$, the adversary prevents the fair classifier from reducing the bias in its predictions w.r.t. the complete data. Moreover, AMMCB requires knowledge of the hypothesis spaces $\Theta_{C}$ and $\Theta_{I}$, which might not be accessible by the adversary in practice. It is still convenient to formalize AMMCB because it defines the ultimate goal of the adversary. However, instead of solving this problem, we focus on the following more tractable surrogate.
@@
Adversarial Missingness for Minimizing the Bias in the Data (AMMBD): Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with a feature matrix $\mathcal{X}$, a hypothesis space $\Theta_{I}$, and a training dataset $\mathcal{V}_{L}$, select $\mathcal{V}_{S}\subset\mathcal{V}$ such that $|\mathcal{V}_{S}|=k$ as to minimize the dataset bias generated by an imputation model $f_{I}^{*}\in\Theta$ over $\mathcal{V}_{L}$.
AMMBD does not depend on $\Theta_{C}$ and its objective is constrained to the training data. The relationship between AMMBD and AMMCB is due to generalization—i.e., the expectation that the bias in $\mathcal{V}_{L}$ will generalize to the predictions from $f_{C}$. However, we show that even AMMDB is NP-hard.

### v5

@@ Abstract
Graph Neural Networks (GNNs) have achieved state-of-the-art results in many relevant tasks where decisions might disproportionately impact specific communities. However, existing work on fair GNNs often assumes that either sensitive attributes are fully observed or they are missing completely at random. We show that an adversarial missingness process can inadvertently disguise a fair model through the imputation, leading the model to overestimate the fairness of its predictions. We address this challenge by proposing Better Fair than Sorry (BFtS), a fair missing data imputation model for sensitive attributes. The key principle behind BFtS is that imputations should approximate the worst-case scenario for fairness—i.e., when optimizing fairness is the hardest. We implement this idea using a 3-player adversarial scheme where two adversaries collaborate against a GNN classifier, and the classifier minimizes the maximum bias. Experiments using synthetic and real datasets show that BFtS often achieves a better fairness $\times$ accuracy trade-off than existing alternatives under an adversarial missingness process.
@@ 1 Introduction
With the increasing popularity of machine learning in high-stakes decision-making, it has become a consensus that these models carry implicit biases that should be addressed to improve the fairness of algorithmic decisions (Ghallab, 2019). The disparate treatment of such models towards African Americans and women has been illustrated in the well-documented COMPAS (Angwin et al., 2022) and Apple credit card (Vigdor, 2019) cases, respectively. While there has been extensive research on fair ML, the proposed solutions have mostly disregarded important challenges that arise in real-world settings. For instance, in many applications, data can be naturally modeled as graphs (or networks) (Dong et al., 2023), representing different objects, their relationships, and attributes, instead of as sequences or images. Moreover, fair ML is also prone to missing data (Little & Rubin, 2019). This is particularly critical because fair algorithms often require knowledge of sensitive attributes that are more likely to be missing due to biases in the collection process or privacy concerns. For instance, census and health surveys exhibit missingness correlated with gender, age, and race (O’Hare & O’Hare, 2019; Weber et al., 2021). Networked data, such as disease transmission studies, face similar issues (Ghani et al., 1998). As a consequence, fair ML methods often rely on missing data imputation, which can introduce errors that compromise fairness (Mansoor et al., 2022; Jeong et al., 2022).
This work investigates the impact of adversarial sensitive value missingness processes on fairness, where the pattern of missing sensitive data is structured to obscure true disparities. If the predictions from the imputation method fail to capture the true distribution of a sensitive attribute, any fairness-aware model trained on the imputed data is prone to inheriting the underlying true bias. More specifically, a key challenge arises when an adversarial missingness process makes the imputed dataset appear to be fair. In graphs, an adversary can exploit the graph structure to manipulate the imputed values and, as a consequence, the fairness-aware model. Prior work on the fairness of graph machine learning assumes sensitive values are Missing Completely At Random (MCAR). However, this assumption rarely holds in practice (O’Hare & O’Hare, 2019; Weber et al., 2021; Ghani et al., 1998; Jeong et al., 2022). Luh (2022) and Fukuchi et al. (2020) provide motivational examples of how fairness can be manipulated by adversarial data collection. Methods that overlook the adversarial missingness process risk producing misleading fairness guarantees w.r.t. the complete data.
@@
Figure 1: In Fig. (a), a graph machine learning algorithm is applied to decide who receives credit (positive or negative) based on a possibly missing sensitive attribute, gender (and binary only for illustrative purposes). As shown in Fig. (b), traditional missing data imputation does not account for outcomes (positive/negative), and thus, their imputed values can under-represent the bias of the complete dataset—demographic parity (DP) is $0.09$ in this example (DP and bias are inversely related). This paper proposes BFtS, an imputation method for graph data that optimizes fairness in the worst-case imputation scenario using adversarial learning, as shown in Fig. (c), where DP is $0.47$.
@@
We summarize our contributions as follows: (1) We investigate theoretically and empirically the potential of an adversarial missingness process to bias a fair GNN; (2) We propose Better Fair than Sorry (BFtS), a novel 3-player adversarial learning framework for the imputation of missing sensitive data that produces worst-case imputed values for fair GNNs and is effective under adversarial missingness processes and even when sensitive attribute information is completely unavailable; and (3) We show empirically that BFtS achieves a better fairness $\times$ accuracy trade-off than the baselines.
@@ 1.1 Related Work
Fairness in graph machine learning. Our work is focused on group fairness Pessach & Shmueli (2022); Dong et al. (2023). Existing work can be grouped into pre-processing, extended objectives, and adversarial learning. Pre-processing methods such as FairOT, FairDrop, FairSIN, EDITS, and Graphair remove bias from the graph before training (Laclau et al., 2021; Spinelli et al., 2021; Yang et al., 2024; Dong et al., 2022; Ling et al., 2023). Objective-based methods—including Fairwalk, Crosswalk, Debayes, MONET, NIFTY, FairVGNN, PFR-AX, FairSAD, and FairGAE—modify GNN losses to learn fair representations (Rahman et al., 2019; Khajehnejad et al., 2022; Buyl & De Bie, 2020; Palowitch & Perozzi, 2019; Agarwal et al., 2021; Wang et al., 2022; Merchant & Castillo, 2023; Zhu et al., 2024; Fan et al., 2021). Adversarial methods such as CFC, FLIP, DKGE, and Debias jointly train GNNs with adversaries for fair prediction (Bose & Hamilton, 2019; Masrour et al., 2020; Arduini et al., 2020; Zhang et al., 2018). All these require fully observed sensitive attributes.
@@
Missing Data Imputation and fairness. Missing data can be imputed using unconditional mean, reconstruction, preferential attachment, autoencoders, etc. (Donders et al., 2006; Pereira et al., 2020; Huisman, 2009). The traditional procedure for handling missing data is independent imputation—i.e., to first impute the missing values and then solve the task (Rossi et al., 2022; Buck, 1960). SAT is a matching-based GNN for graphs with missing attributes (Chen et al., 2020), but it does not account for fairness. Training a classifier from imputed data can amplify the bias of a machine learning model, as discussed in (Subramonian et al., 2022; Zhang & Long, 2021; Feng et al., 2023; Guha et al., 2024; Martínez-Plumed et al., 2019; Fricke, 2020; Mansoor et al., 2022). Some studies try to generate fair feature imputations (Subramonian et al., 2022; Feng et al., 2023; Jeong et al., 2022; Zhang & Long, 2022). However, as the approaches discussed earlier, these methods do not consider missing sensitive information and require full knowledge of the sensitive attributes for fairness intervention.
@@
Fairness with missing sensitive attributes. One extreme assumption is the complete unavailability of sensitive information. For example, Hashimoto et al. (2018) considers the worst-case distribution over group sizes, Lahoti et al. (2020) reweighs training samples adversarially, Chai & Wang (2022) minimizes a top-$k$ average loss, and Zhao et al. (2022) reduces correlation between predictions and features associated with sensitive attributes. Other approaches include class re-balancing (Yan et al., 2020) and soft labels from an overfitted teacher (Chai et al., 2022). In practice, partial sensitive information is often available and improves fairness (Chai & Wang, 2022), but these models fail to leverage it. FairGNN (Dai & Wang, 2021) assumes limited sensitive data and imputes missing values independently, FairAC (Guo et al., 2023) uses only observed attributes without imputation, and RNF (Du et al., 2021) generates proxy annotations using generalized cross-entropy (Zhang & Sabuncu, 2018). Similar to RNF, our approach handles completely or partially missing sensitive information. BFtS applies a 3-player scheme to minimize the maximum possible bias and outperforms FairGNN and RNF in terms of fairness $\times$ accuracy. Minimizing the maximum instead of the average risk, similar to our method, has been shown to achieve better guarantees (Shalev-Shwartz & Wexler, 2016). A similar minmax approach has been applied to maximize the robustness and accuracy of uncertainty models (Löfberg, 2003; Lanckriet et al., 2002; Chen et al., 2017; Fauß et al., 2021). In (Nguyen et al., 2017; Vandenhende et al., 2019), a 3-player adversarial network is proposed to improve the classification and stability of adversarial learning.
Adversarial attacks on fairness. Recent work has investigated data poisoning as a means to adversarially degrade model fairness (Mehrabi et al., 2021; Solans et al., 2020). UnfairTrojan and TrojFair introduce backdoor attacks specifically aimed at reducing model fairness (Furth et al., 2024; Zheng et al., 2023). In the context of graph neural networks (GNNs), most adversarial fairness attacks modify graph topology: Hussain et al. (2022); Zhang et al. (2024) perturb edges, NIFA injects nodes via uncertainty maximization and homophily enhancement (Luo et al., 2024), and FATE applies bilevel meta-learning for poisoning (Kang et al., 2024). Thus, prior work has primarily focused on poisoning, node injection, or structural perturbations, while attacks that compromise fairness solely by manipulating sensitive-attribute missingness remain unexplored. In this study, we focus on adversarial missingness rather than data poisoning, as it represents a more practical and plausible threat model in many real-world systems. Prior work on adversarial missingness in causal structure learning (Koyuncu et al., 2023; 2024) highlights that when data authenticity is enforced (e.g., via cryptographically signed sensor records), an adversary cannot modify values or inject fabricated samples without violating digital-signature constraints. Such tampering would introduce inconsistencies across logs, backups, and signatures, making it infeasible under standard integrity guarantees. In contrast, selectively withholding, dropping, or failing to log certain fields does not break these integrity checks and is indistinguishable from common pipeline failures. Under this threat model, the adversary is therefore limited to partially concealing existing data and can strategically choose which sensitive attributes to withhold. This makes adversarial missingness a plausible and practically relevant mechanism for undermining fairness.
@@ 2 Preliminaries
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X},\mathcal{S})$ be an undirected graph where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, $\mathcal{X}$ are node attributes, and $\mathcal{S}$ is the set of sensitive attributes. The matrix $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix of $\mathcal{G}$ where $A_{uv}=1$ if there is an edge between $u$ and $v$ and $A_{uv}=0$, otherwise. We focus on the setting where sensitive attributes might be missing, and only $\mathcal{V}_{S}\subseteq\mathcal{V}$ nodes include the information of the sensitive attribute $s_{v}$ for a node $v$. The sensitive attribute forms two groups, which are often called sensitive ($s_{v}=1$) and non-sensitive ($s_{v}=0$).
While our work can be generalized to other fairness-aware tasks, we will focus on binary fair node classification, where the goal is to learn a classifier $f_{C}$ to predict node labels $y_{v}\in\{0,1\}$ based on a training set $\mathcal{V}_{L}\subseteq\mathcal{V}$. Without loss of generality, we assume that the class $y=1$ is the desired one (e.g., receive credit or bail). Given a classification loss $\mathcal{L}_{class}$ and a group fairness loss $\mathcal{L}_{bias}$, the goal is to learn the parameters $\theta_{class}$ of $f_{class}$ by minimizing their combination:
$\theta_{class}^{*}=\arg\min_{\theta_{class}}\mathcal{L}_{class}+\alpha\mathcal{L}_{bias}$
where $\mathcal{L}_{bias}$ measures the impact of the sensitive attribute $s_{v}$ over the predictions from $f_{class}$ and $\alpha$ is a hyperparameter. The main challenge addressed in this paper is how to handle missing sensitive attributes. In this scenario, one can apply an independent imputation model $s_{v}\approx f_{imp}(v)$ before training the fair classifier. However, the fairness of the resulting classifier will be highly dependent on the accuracy of $f_{imp}$, as will be discussed in more detail in the next section.
@@ 3 Introducing Bias via an Adversarial Missingness Process
We investigate the impact of missing sensitive values on fairness in graph machine learning. Datasets with substantial missingness typically require imputation. In this section, we focus on how the missingness process (i.e., the process that generates missing values) can lead to (intended or unintended) biases in a fair model trained using the imputed data, leading fair models to underestimate bias and remain unfair relative to the complete data.
We can describe the missingness process using a threat model. The asset is the graph $\mathcal{G(\mathcal{V},\mathcal{E},\mathcal{X},\mathcal{S})}$, and the threat is increasing the bias of a node classification model $f_{class}$ applied to $\mathcal{G}$. The vulnerability arises because the adversary can select a subset of sensitive attributes $\mathcal{V}\setminus\mathcal{V}_{S}$ to be missing, thereby inducing a strategically biased missingness pattern. Our mitigation combines missing-data imputation and fair node classification to counter such adversarial missingness. The attacker’s capabilities and constraints are described next in the adversarial missingness formulations introduced in this section.
The simplest missingness process for sensitive values is missing completely at random—i.e., the probability of a value being missing is independent of the data. However, in practical scenarios, missing values are rarely random (O’Hare & O’Hare, 2019; Weber et al., 2021; Ghani et al., 1998). We focus on the particular case where the missingness process is adversarial. We prove that manipulating the missingness process to induce bias optimally is computationally hard. However, a simple heuristic can effectively introduce bias to an independent imputation model that simply tries to maximize imputation accuracy. We formulate the problem for a given number of observed sensitive values $k$.
@@
Adversarial Missingness Against Fair Classification (AMAFC): Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, select a set of nodes $\mathcal{V}_{S}^{*}$ that maximizes the bias of a fair classifier with parameters $\theta_{class}^{*}$ trained with an imputation model with parameters $\theta_{imp}^{*}$ trained with $\mathcal{V}_{S}$:
$\mathcal{V}_{S}^{*}$ $={\arg\max}_{\mathcal{V}_{S}\in\mathcal{V},|\mathcal{V}_{S}|=k}\mathcal{L}_{bias}(\theta_{class}^{*},\mathcal{V})$
$\quad\text{s.t.}\quad\theta_{class}^{*}={\arg\min}_{\theta_{class}}\mathcal{L}_{class}(\theta_{class})+\alpha\mathcal{L}_{bias}(\theta_{class},\theta_{imp}^{*},\mathcal{V}_{S})$
$\quad\quad\text{s.t.}\quad\theta_{imp}^{*}={\arg\min}_{\theta_{imp}}\mathcal{L}_{imp}(\theta_{imp},\mathcal{V}_{S})$
where we assume that the adversary can compute the bias $\mathcal{L}_{bias}(\theta_{class}^{*},\mathcal{V})$ based on all sensitive attributes $s_{v}$, while the classifier can only estimate its bias based on a combination of nodes $\mathcal{V}_{S}$ with observed $s_{v}$ and imputed values produced with $\theta_{imp}^{*}$. While AMAFC describes the objective of an idealized adversary, it is impractical due to its associated complexity (tri-level optimization).
@@
Adversarial Missingness against Data Bias (AMADB): Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, select a set of nodes $\mathcal{V}_{S}^{*}$ that minimizes the bias in the labels estimated using an imputation model with parameters $\theta_{imp}^{*}$ trained with $\mathcal{V}_{S}$:
$\mathcal{V}_{S}^{*}$ $={\arg\min}_{\mathcal{V}_{S}\in\mathcal{V},|\mathcal{V}_{S}|=k}\mathcal{L}_{bias}(\theta_{imp}^{*},\mathcal{V}_{S},\mathcal{V}_{L})$
$\quad\text{s.t.}\quad\theta_{imp}^{*}={\arg\min}_{\theta_{imp}}\mathcal{L}_{imp}(\theta_{imp},\mathcal{V}_{S})$
where $\mathcal{L}_{bias}(\theta_{imp}^{*},\mathcal{V}_{S},\mathcal{V}_{L})$ is computed based on labels instead of $f_{class}$. By minimizing the bias computed based on imputed values and labels, AMADB attempts to misguide any classifier that relies on such imputation to mitigate bias. AMADB is more tractable than AMFC, but still NP-hard.